# MixAT: Combining Continuous and Discrete Adversarial Training for LLMs

### v1

@@ 4.2 Main Results
In Table 1 we compare MixAT with stat-of-the-art adversarial training methods (R2D2 [5], CAT [7], LAT [9]) and prompt-based defences (RPO [20], PAT [19]), as well as our baseline PAP-AT and the variant DualAT. We show that MixAT achieves the best tradeoff between robustness and utility, outperforming other methods on both metrics.
@@
Finally, we show that MixAT is also effective on larger models, achieving very low ALO-ASR on Qwen2.5-14B and Qwen2.5-32B, while maintaining competitive utility scores, as well as on Mistral-7B [33] and Llama3.1-8B [25] in Section B.7.
@@ 4.3 MixAT Ablation Studies
(a) Effect of $\alpha_{PAP}$ on model robustness.
(b) MixAT vs DualAT ALO-ASR [%] $\downarrow$.
@@
Continuous vs Discrete trade-off In Fig. 3 we examine the effect of varying the amount ($\alpha\in[0,1]$) of mixed attack samples used throughout the training on model robustness. Expectedly, we observe that for low $\alpha$ values, the models are less robust against PAP attacks. On the other hand, models trained with high $\alpha$ values can become less robust against Direct Requests because the model does not see enough clean malicious samples during training. We choose $\alpha=0.5$ as the default value for our main experiments as a balanced trade-off between direct and paraphrased samples, since we also observe more stability and robustness in this region. We present the full results (including utilities) in Table 3, and observe that the $\alpha$ ratio does not seem to have a significant effect on model utility.
@@
MixAT vs DualAT We further examine the robustness differences between MixAT and DualAT in Fig. 3. While MixAT trains the model using combined attacks, DualAT separately trains the model on continuous and discrete attacks. When $\alpha=0$, both methods are equivalent to CAT training, while for $\alpha=1$ only DualAT corresponds to PAP-AT training. We observe that MixAT generally outperforms the simpler DualAT, indicat ing that training with the our attack combination is more effective than directly combining the continuous and discrete training.
@@
Incorporating GCG samples into MixAT training Our results show that while MixAT achieves almost perfect robustness against most prompt-based attacks, it is still slightly more vulnerable to GCG attacks. To further enhance model’s robustness against GCG attacks, we experiment with including some GCG samples in the training process. To do that we adapt our training mix to include $\alpha_{GCG+C}=10\%$ GCG samples with continuous attacks on top, $\alpha_{PAP+C}=45\%$ PAP samples with continuous attacks on top, and $\alpha_{C}=45\%$ clean samples with continuous attacks on top. In Table 1 we show that this approach improves the robustness against GCG attacks, while keeping the robustness against other attacks and the utility scores similar to the default MixAT training. However, even by adding only 10% GCG attacks and only running them for 100 steps (as opposed to 500 steps used for attacking), we observed a 5 times increase in training time when compared to the base MixAT (Table 2). While the results obtained by mixing paraphrasing, continuous and suffix attack samples are already promising, this method could benefit from further exploration to obtain better hyperparameters and reduced computational costs.
@@
MixAT static training To validate the importance of using dynamically generated attacks while training, we also experiment with a static version of MixAT, where all adversarial samples used for training are generated on the base model. We observe that this approach leads to a significant drop in robustness, with the ALO-ASR increasing from 12.5% to 25% on Zephyr-7B and from 25% to 52.5% on Llama3-8B (for detailed results see Table 12 in Appendix B). This indicates that the dynamic generation of adversarial samples during training is crucial for achieving high robustness.
4.4 Scaling the LoRA weights
One of the main challenges of adversarial training is to find the optimal trade-off between adversarial robustness and utility. Intuitively, we can regulate the effect of LoRA adapters [34] by scaling their weights with a constant $\lambda$. This approach may yield different trade-offs depending on the chosen $\lambda$ values. To test our hypothesis, we evaluate LoRA-scaled variants of the MixAT and CAT Zephyr-7B models on utility and adversarial robustness benchmarks. This involves multiplying the low-rank matrix $A$ by the constant $\lambda$: $W=W_{0}+(\lambda A)B$, where $W_{0}$ denotes the original weight matrix, A and B are low-rank matrices used to introduce perturbations to the model weights, and $\lambda$ is the scaling factor. We scaled $\lambda$ from 0.0 to 1.5 with a step of 0.25. As shown in Fig. 5, the ASR (Adversarial Success Rate) decreases as the magnitude of LoRA scaling increases, supporting our hypothesis about regulating the effect of LoRA. This decline in ASR is more rapid for the MixAT model compared to the CAT model, indicating the higher effectiveness of the MixAT training.
On the other hand, increasing the strength of the LoRA adapter also gradually reduces the models’ utility, confirming the inherent trade-off between utility and robustness. However, we observe that utility scores of MixAT are consistently similar or higher than those of CAT across all $\lambda$. In general, for $\lambda>1$, the models begin to significantly lose utility through over-refusal. In the range $10<\lambda<$ 1, utility is not significantly degraded (even slightly improving on some tasks). For the MixAT, we observe that values of $\lambda$ between 0.5 and 1.0 yield good robustness with minimal utility losses.
@@ 4.5 Effect of Temperature on Robustness Evaluation
In Fig. 5 we examine the effect of temperature on the robustness of the LLama-3-8B model trained with MixAT. For each temperature value, we sample the model 10 times for each harmful prompt. We analyze the distribution of harmful generations across temperatures, reporting the percentage of prompts for which the model produces at least one harmful response (1/10 ASR), all harmful responses (10/10) and the ASR averaged across all samples (Avg. ASR). We observe that the average ASR score does not change significantly with temperature, but the likelihood of generating at least one harmful response in multiple tries increases. This is consistent with the findings of Raina et al. [35], who show that the model’s robustness is greatly correlated to the first few generated tokens.
@@ 4.6 Further Experiments on Robustness (Evaluations)
We also conduct a series of additional experiments to further investigate the robustness and utility of our models. We present the key findings here, with extended details and data provided in Appendix B.
Discussion on Randomness The training and evaluation of LLMs are inherently random due to multiple reasons. In Section B.2 we investigate the impact of randomness in the training process (Table 5), and compare it with the randomness in the evaluation process (Table 4). We observe that models that are very robust (low ASR scores) or very unrobust (high ASR scores) are less affected by randomness, while models with intermediate ASR scores are more affected.
@@
Trained Model GPUs VRAM Train Train Total Est.
@@
Impact of Model Quantization Model quantization is a frequently used technique to decrease memory costs. We examine the effects of quantization on robustness during training and evaluation in Section B.3. We observe that quantization slightly improves the robustness, but also lowers the capabilities of the model. This is valid when quantization is applied both during training and during evaluation.
Evaluating on more samples In Section B.4 we evaluate the robustness of some of our models on the whole HarmBench test set, as well as a subset of the XSTest-Harmful set. We find the robustness trends with respect to models and attacks to be mostly stable in our experiments. This indicates that evaluating on the 40-sample subset is a good proxy for quantifying models’ robustness.
Transferability of attacks Prior work [4] has investigated the transferability of adversarial attacks across regular LLM models. In Section B.5, we assess how transferability is affected when adversarially trained models are used. We observe that transferability in this scenario is harder.
@@ 4.7 Training Time and Costs
Next, we report the training times of the main methods examined in this work in Table 2. As shown there, we train all of the models using either NVIDIA A100-40GB or NVIDIA H200 GPUs. The results show that the compute time and resources required for MixAT are lower than those required for CAT and R2D2, and only slightly higher than those required for LAT. We observe that the costs scale roughly linearly with the model size. This indicates that our method is efficient and can be applied to larger models without significant computational overhead.
@@ 5 Limitations
The main limitations of our work lie in the evaluation process, which, while improved over prior methods, continues to pose significant challenges. As discussed in Section 4, randomness during both training and evaluation introduces considerable variability. This issue is compounded by the high computational cost of evaluation, which forces us to use a limited number of samples and leads to higher variance. Additionally, the ambiguity in determining the malicious intent of certain evaluation samples in datasets like MT-Bench further contributes to noise in the reported metrics.
@@ B.1 Extended Ablation Results
Method $\alpha_{D}$ Utility Benchmarks [%] Attack Success Rate [%]
@@
CAT short 0.0 81.5 54.9 56.2 95.0 57.4 45.6 0.0 47.5 30.0 25.0 0.0 12.5 2.5 75.0
MixAT 0.1 81.5 55.1 56.2 95.0 56.4 57.6 2.5 12.5 10.0 2.5 2.5 7.5 5.0 22.5
MixAT 0.2 81.6 55.4 56.0 97.5 53.2 52.4 2.5 7.5 10.0 2.5 0.0 5.0 2.5 20.0
MixAT 0.3 81.2 54.0 56.3 85.0 54.3 41.2 2.5 2.5 5.0 2.5 2.5 2.5 2.5 10.0
MixAT 0.4 81.7 54.0 56.0 97.5 54.0 59.2 0.0 5.0 0.0 0.0 0.0 10.0 0.0 12.5
MixAT 0.5 81.4 54.0 55.8 97.5 54.3 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
MixAT 0.6 81.4 54.3 56.0 95.0 53.6 45.6 0.0 0.0 2.5 2.5 0.0 2.5 0.0 5.0
MixAT 0.7 81.7 55.8 56.1 100.0 53.5 62.8 2.5 0.0 10.0 2.5 0.0 15.0 5.0 20.0
MixAT 0.8 81.4 54.1 56.1 97.5 55.8 86.4 7.5 2.5 2.5 0.0 0.0 5.0 5.0 12.5
MixAT 0.9 81.6 54.6 56.1 87.5 50.5 53.6 0.0 0.0 0.0 0.0 0.0 5.0 2.5 7.5
MixAT 1.0 81.9 54.8 56.3 97.5 55.8 93.6 7.5 0.0 0.0 0.0 0.0 2.5 10.0 12.5
DualAT 0.1 81.6 54.9 56.3 95.0 54.4 51.2 5.0 17.5 17.5 35.0 2.5 15.0 7.5 47.5
DualAT 0.2 82.1 55.1 56.2 92.5 53.1 48.0 2.5 10.0 37.5 60.0 0.0 10.0 5.0 72.5
DualAT 0.3 81.6 54.5 56.0 92.5 55.3 62.0 0.0 15.0 22.5 50.0 0.0 22.5 5.0 65.0
DualAT 0.4 82.1 54.9 56.0 95.0 54.1 51.2 2.5 10.0 12.5 22.5 2.5 20.0 2.5 37.5
DualAT 0.5 81.8 54.4 56.1 85.0 54.2 47.2 2.5 2.5 10.0 15.0 0.0 10.0 2.5 22.5
DualAT 0.6 81.9 55.0 56.1 92.5 53.4 49.6 2.5 2.5 12.5 12.5 0.0 15.0 2.5 27.5
DualAT 0.7 82.2 54.0 56.0 95.0 49.5 49.6 2.5 0.0 2.5 7.5 0.0 15.0 2.5 15.0
DualAT 0.8 81.4 53.9 55.8 97.5 52.7 60.4 2.5 0.0 0.0 2.5 0.0 20.0 7.5 20.0
DualAT 0.9 81.0 54.6 55.7 97.5 51.1 60.8 2.5 2.5 2.5 0.0 0.0 30.0 2.5 32.5
PAP-AT 1.0 82.2 54.5 56.0 100.0 54.1 94.0 17.5 2.5 5.0 15.0 2.5 55.0 57.5 77.5
Zephyr-7B Base 81.0 55.2 56.2 100.0 61.4 98.8 85.0 87.5 85.0 97.5 90.0 85.0 100.0 100.0
In Table 3 we present full results of MixAT (PAP+c & CA) and DualAT (PAP & CA) style trainings on Zephyr-7B model with different $\alpha$ ratio of PAP examples utilized throughout the training. The results suggest that models trained with MixAT are generally more robust than those trained with DualAT (with the examination of ALO scores corresponding to different $\alpha$ percentages). It can be also examined that low ALO (better robustness) often leads to worse performance on MT-Bench.

### v2

@@ 4.2 Main Results
In Table˜1 we compare MixAT with stat-of-the-art adversarial training methods (R2D2 [5], CAT [7], LAT [9]) and prompt-based defences (RPO [21], PAT [20]), as well as our baseline PAP-AT and the variant DualAT. We show that MixAT achieves the best tradeoff between robustness and utility, outperforming other methods on both metrics.
(a) Effect of $\alpha_{PAP}$ on model robustness.
(b) MixAT vs DualAT ALO-ASR [%] $\downarrow$.
@@
Finally, we show that MixAT is also effective on larger models, achieving very low ALO-ASR on Qwen2.5-14B and Qwen2.5-32B, while maintaining competitive utility scores, as well as on Mistral-7B [36] and Llama3.1-8B [28] in Section˜B.7.
@@ 4.3 MixAT Ablation Studies
Continuous vs Discrete trade-off In Fig.˜3 we examine the effect of varying the amount ($\alpha\in[0,1]$) of mixed attack samples used throughout the training on model robustness. Expectedly, we observe that for low $\alpha$ values, the models are less robust against PAP attacks. On the other hand, models trained with high $\alpha$ values can become less robust against Direct Requests because the model does not see enough clean malicious samples during training. We choose $\alpha=0.5$ as the default value for our main experiments as a balanced trade-off between direct and paraphrased samples, since we also observe more stability and robustness in this region. We present the full results (including utilities) in Table˜3, and observe that the $\alpha$ ratio does not seem to have a significant effect on model utility.
@@
MixAT vs DualAT We further examine the robustness differences between MixAT and DualAT in Fig.˜3. While MixAT trains the model using combined attacks, DualAT separately trains the model on continuous and discrete attacks. When $\alpha=0$, both methods are equivalent to CAT training, while for $\alpha=1$ only DualAT corresponds to PAP-AT training. We observe that MixAT generally outperforms the simpler DualAT, indicat ing that training with the our attack combination is more effective than directly combining the continuous and discrete training.
@@
Incorporating GCG samples into MixAT training Our results show that while MixAT achieves almost perfect robustness against most prompt-based attacks, it is still slightly more vulnerable to GCG attacks. To further enhance model’s robustness against GCG attacks, we experiment with including some GCG samples in the training process. To do that we adapt our training mix to include $\alpha_{GCG+C}=10\%$ GCG samples with continuous attacks on top, $\alpha_{PAP+C}=45\%$ PAP samples with continuous attacks on top, and $\alpha_{C}=45\%$ clean samples with continuous attacks on top. In Table˜1 we show that this approach improves the robustness against GCG attacks, while keeping the robustness against other attacks and the utility scores similar to the default MixAT training. However, even by adding only 10% GCG attacks and only running them for 100 steps (as opposed to 500 steps used for attacking), we observed a 5 times increase in training time when compared to the base MixAT (Table˜2). While the results obtained by mixing paraphrasing, continuous and suffix attack samples are already promising, this method could benefit from further exploration to obtain better hyperparameters and reduced computational costs.
@@
MixAT static training To validate the importance of using dynamically generated attacks while training, we also experiment with a static version of MixAT, where all adversarial samples used for training are generated on the base model. We observe that this approach leads to a significant drop in robustness, with the ALO-ASR increasing from 12.5% to 25% on Zephyr-7B and from 25% to 52.5% on Llama3-8B (for detailed results see Table˜12 in Appendix˜B). This indicates that the dynamic generation of adversarial samples during training is crucial for achieving high robustness.
@@
4.4 Scaling the LoRA weights
One of the main challenges of adversarial training is to find the optimal trade-off between adversarial robustness and utility. Intuitively, we can regulate the effect of LoRA adapters [37] by scaling their weights with a constant $\lambda$. This approach may yield different trade-offs depending on the chosen $\lambda$ values. To test our hypothesis, we evaluate LoRA-scaled variants of the MixAT and CAT Zephyr-7B models on utility and adversarial robustness benchmarks. This involves multiplying the low-rank matrix $A$ by the constant $\lambda$: $W=W_{0}+(\lambda A)B$, where $W_{0}$ denotes the original weight matrix, A and B are low-rank matrices used to introduce perturbations to the model weights, and $\lambda$ is the scaling factor. We scaled $\lambda$ from 0.0 to 1.5 with a step of 0.25. As shown in Fig.˜5, the ASR (Adversarial Success Rate) decreases as the magnitude of LoRA scaling increases, supporting our hypothesis about regulating the effect of LoRA. This decline in ASR is more rapid for the MixAT model compared to the CAT model, indicating the higher effectiveness of the MixAT training.
On the other hand, increasing the strength of the LoRA adapter also gradually reduces the models’ utility, confirming the inherent trade-off between utility and robustness. However, we observe that utility scores of MixAT are consistently similar or higher than those of CAT across all $\lambda$. In general, for $\lambda>1$, the models begin to significantly lose utility through over-refusal. In the range $0<\lambda<1$, utility is not significantly degraded (even slightly improving on some tasks). For the MixAT, we observe that values of $\lambda$ between 0.5 and 1.0 yield good robustness with minimal utility losses.
@@ 4.5 Effect of Temperature on Robustness Evaluation
In Fig.˜5 we examine the effect of temperature on the robustness of the LLama-3-8B model trained with MixAT. For each temperature value, we sample the model 10 times for each harmful prompt. We analyze the distribution of harmful generations across temperatures, reporting the percentage of prompts for which the model produces at least one harmful response (1/10 ASR), all harmful responses (10/10) and the ASR averaged across all samples (Avg. ASR). We observe that the average ASR score does not change significantly with temperature, but the likelihood of generating at least one harmful response in multiple tries increases. This is consistent with the findings of Raina et al. [38], who show that the model’s robustness is greatly correlated to the first few generated tokens.
@@ 4.6 Further Experiments on Robustness (Evaluations)
We also conduct a series of additional experiments to further investigate the robustness and utility of our models. We present the key findings here, with extended details and data provided in Appendix˜B.
Discussion on Randomness The training and evaluation of LLMs are inherently random due to multiple reasons. In Section˜B.2 we investigate the impact of randomness in the training process (Table˜5), and compare it with the randomness in the evaluation process (Table˜4). We observe that models that are very robust (low ASR scores) or very unrobust (high ASR scores) are less affected by randomness, while models with intermediate ASR scores are more affected.
@@
Trained
Model
GPUs VRAM Train Train Total Est.
@@
Impact of Model Quantization Model quantization is a frequently used technique to decrease memory costs. We examine the effects of quantization on robustness during training and evaluation in Section˜B.3. We observe that quantization slightly improves the robustness, but also lowers the capabilities of the model. This is valid when quantization is applied both during training and during evaluation.
Evaluating on more samples In Section˜B.4 we evaluate the robustness of some of our models on the whole HarmBench test set, as well as a subset of the XSTest-Harmful set. We find the robustness trends with respect to models and attacks to be mostly stable in our experiments. This indicates that evaluating on the 40-sample subset is a good proxy for quantifying models’ robustness.
Transferability of attacks Prior work [4] has investigated the transferability of adversarial attacks across regular LLM models. In Section˜B.5, we assess how transferability is affected when adversarially trained models are used. We observe that transferability in this scenario is harder.
@@ 4.7 Training Time and Costs
Next, we report the training times of the main methods examined in this work in Table˜2. As shown there, we train all of the models using either NVIDIA A100-40GB or NVIDIA H200 GPUs. The results show that the compute time and resources required for MixAT are lower than those required for CAT and R2D2, and only slightly higher than those required for LAT. We observe that the costs scale roughly linearly with the model size. This indicates that our method is efficient and can be applied to larger models without significant computational overhead.
@@ 5 Limitations
The main limitations of our work lie in the evaluation process, which, while improved over prior methods, continues to pose significant challenges. As discussed in Section˜4, randomness during both training and evaluation introduces considerable variability. This issue is compounded by the high computational cost of evaluation, which forces us to use a limited number of samples and leads to higher variance. Additionally, the ambiguity in determining the malicious intent of certain evaluation samples in datasets like MT-Bench further contributes to noise in the reported metrics.
6 Discussion
In this section, we provide a broader discussion of the role of adversarial training in LLM safety, as well as the current state of adversarial LLM evaluation.
Can Adversarial Training Fully Mitigate Malicious Requests? While we consider adversarial training a foundational tool for defending LLMs against malicious inputs, it is unlikely to fully resolve the problem of harmful requests. In particular, we find that with many adversarial training techniques, LLMs can be significantly strengthened to resist specific (known) attack types. However, in contrast to conventional neural network classifiers, the input-output space of LLMs is substantially broader, with adversarially trained models often remaining vulnerable to adaptive attackers who craft novel strategies beyond the training space. Moreover, adversarial training alone does not address the deeper ethical and contextual understanding that is critical for appropriately responding to harmful inputs—and, in some cases, even for categorizing them as harmful. A holistic defense strategy must therefore combine adversarial training with complementary methods, such as enhanced filtering systems, context-aware response generation, and ongoing model evaluation and refinement.
Challenges in Evaluation: Are Current Datasets Adequate? Any statement about the robustness of an (adversarially trained) LLM is fundamentally tied to the quality of the evaluation datasets used. During our evaluation, we found that many existing datasets exhibit specific flaws that prevent more holistic statements about LLM robustness. In particular, while these datasets collectively address various attack types, their scope is typically insufficient to test resilience against more powerful adversaries. Importantly, overemphasizing a single attack type can lead to miscontextualized results claiming high overall robustness despite failures under other, equally practical attacks. Additionally, many datasets disproportionately emphasize certain attack categories, such as toxicity, while underrepresenting other forms of harmful or manipulative input. Given the lack of an apparent “best attack,” we consider our proposed ALO-ASR a practical post hoc improvement over prior approaches. Nevertheless, to better evaluate and improve LLM robustness, there remains a pressing need for richer, more diverse datasets encompassing a broader range of adversarial examples, including dynamic and context-sensitive attacks.
@@ Acknowledgments
Part of this work has been done under the SERI grant SAFEAI (Certified Safe, Fair and Robust Artificial Intelligence, contract no. MB22.00088). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or European Commission. Neither the European Union nor the European Commission can be held responsible for them. The work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) (SERI-funded ERC Consolidator Grant).
@@ B.1 Extended Ablation Results
Method
$\alpha_{D}$
Utility Benchmarks [%] Attack Success Rate [%]
@@
CAT short 0.0 81.5 54.9 56.2 95.0 55.4 45.6 0.0 47.5 30.0 25.0 0.0 12.5 2.5 75.0
MixAT 0.1 81.5 55.1 56.2 95.0 55.4 57.6 2.5 12.5 10.0 2.5 2.5 7.5 5.0 22.5
MixAT 0.2 81.6 55.4 56.0 97.5 50.7 52.4 2.5 7.5 10.0 2.5 0.0 5.0 2.5 20.0
MixAT 0.3 81.2 54.0 56.3 85.0 47.6 41.2 2.5 2.5 5.0 2.5 2.5 2.5 2.5 10.0
MixAT 0.4 81.7 54.0 56.0 97.5 50.9 59.2 0.0 5.0 0.0 0.0 0.0 10.0 0.0 12.5
MixAT 0.5 81.4 54.0 55.8 97.5 52.2 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
MixAT 0.6 81.4 54.3 56.0 95.0 49.1 45.6 0.0 0.0 2.5 2.5 0.0 2.5 0.0 5.0
MixAT 0.7 81.7 55.8 56.1 100.0 51.8 62.8 2.5 0.0 10.0 2.5 0.0 15.0 5.0 20.0
MixAT 0.8 81.4 54.1 56.1 97.5 53.1 86.4 7.5 2.5 2.5 0.0 0.0 5.0 5.0 12.5
MixAT 0.9 81.6 54.6 56.1 87.5 47.1 53.6 0.0 0.0 0.0 0.0 0.0 5.0 2.5 7.5
MixAT 1.0 81.9 54.8 56.3 97.5 50.8 93.6 7.5 0.0 0.0 0.0 0.0 2.5 10.0 12.5
DualAT 0.1 81.6 54.9 56.3 95.0 53.3 51.2 5.0 17.5 17.5 35.0 2.5 15.0 7.5 47.5
DualAT 0.2 82.1 55.1 56.2 92.5 51.7 48.0 2.5 10.0 37.5 60.0 0.0 10.0 5.0 72.5
DualAT 0.3 81.6 54.5 56.0 92.5 50.3 62.0 0.0 15.0 22.5 50.0 0.0 22.5 5.0 65.0
DualAT 0.4 82.1 54.9 56.0 95.0 51.8 51.2 2.5 10.0 12.5 22.5 2.5 20.0 2.5 37.5
DualAT 0.5 81.8 54.4 56.1 85.0 51.1 47.2 2.5 2.5 10.0 15.0 0.0 10.0 2.5 22.5
DualAT 0.6 81.9 55.0 56.1 92.5 50.1 49.6 2.5 2.5 12.5 12.5 0.0 15.0 2.5 27.5
DualAT 0.7 82.2 54.0 56.0 95.0 48.8 49.6 2.5 0.0 2.5 7.5 0.0 15.0 2.5 15.0
DualAT 0.8 81.4 53.9 55.8 97.5 51.3 60.4 2.5 0.0 0.0 2.5 0.0 20.0 7.5 20.0
DualAT 0.9 81.0 54.6 55.7 97.5 49.8 60.8 2.5 2.5 2.5 0.0 0.0 30.0 2.5 32.5
PAP-AT 1.0 82.2 54.5 56.0 100.0 52.6 94.0 17.5 2.5 5.0 15.0 2.5 55.0 57.5 77.5
Zephyr-7B Base 81.0 55.2 56.2 100.0 60.3 98.8 85.0 87.5 85.0 97.5 90.0 85.0 100.0 100.0
In Table˜3 we present full results of MixAT (PAP+c & CA) and DualAT (PAP & CA) style trainings on Zephyr-7B model with different $\alpha$ ratio of PAP examples utilized throughout the training. The results suggest that models trained with MixAT are generally more robust than those trained with DualAT (with the examination of ALO scores corresponding to different $\alpha$ percentages). It can be also examined that low ALO (better robustness) often leads to worse performance on MT-Bench.